# Chenjie Zhong    Zhipeng Li†,*    Shangzhi Xu Xiaohu Li

### v1

@@ Abstract
Variational Mode Decomposition (VMD) plays an important role in many scientific areas, especially for the area of signal processing. Unlike the traditional Fourier paradigm, it makes decomposition of a signal possible without any predefined function basis, which gives unprecedented flexibilities while handling narrow-banded signals of varieties. However, determining the number and central frequencies of intrinsic mode functions are still open questions in that few studies has been proposed to give a complete method that can theoretically guarantee the global convergence during decomposition. In this article, we propose a globally convergent numerical optimization method based on variational convex optimization to automatically determine the number and central frequency of IMFs without any prior knowledge for narrow banded signals, as long as the spectra of each sub-band are identifiable. Our method focuses on finding the support baseline of the spectral function, and further separating the significant frequency band regions above the support baseline. Unlike pioneer works that focus on optimizations on complex field, our method achieves obtaining the number of decomposed IMFs and center frequencies in real field by combining variational calculus, convex optimization, and numerical solutions of differential equations in real field and theoretical analysis shows our algorithm is guaranteed terminate to one of the optimum as close as possible. Experiments also shows that our algorithm converges quickly and can be used in practical engineering to determine the prior information of the number of intrinsic modes and evaluate initial center frequency as prior information to continue the VMD procedure.
@@
narrow-banded signals,variational optimization, Intrinsic Mode Function(IMF)
@@ 1 Introduction
Variational Mode Decomposition (VMD) [5] decomposes original signals into multiple Intrinsic Mode Functions (IMF) with limited bandwidth with the help of variational optimization. The core of VMD is to minimize the sum of the bandwidths with respect to the estimated IMFs as well as to ensure the sum of all the IMFs is equal to the original signal. The paradigm of VMD is a significant progress in the field of signal processing in that it not only inherits the core of EMD that breaks the traditional Fourier-transform alike framework for signal decomposition where a signal must be decomposed in terms of fixed signal basis like sine and cosine functions, but is also considered as the first trial to solve the signal decomposition problem with the help of variational optimization. VMD plays an important role in modern signal decomposition after proposition. In the field of mechanics, [14] uses VMD to decompose the vibration signals of planetary gearboxes under variable-speed conditions, in biomedical field, VMD can be used to diagnose the seizure of epileptic by retrieving the feature of decomposed Electroencephalography (EEG) signals [21] and can also help to denoise the Electrocardiogram (ECG) signals [9]. In the field of Earth Physics signals, it not only helps to increase the Signal-to-Noise Ratio(SNR) by combining with adaptive non-local mean filtering [7], but also helps the prediction of monthly runoff by decomposing the signal into trend, periodic and random components as the input feature of downside neural networks [16]. Besides, VMD also provides the potential for the channel identification in communications [1] and helps to the visualization of the protein structure [15]. We show the detail of VMD in Supplementary Materials Section 1. Although VMD has provided flexibility in decomposing signals, there are factors that hinder its further applications in that the number of intrinsic mode function must be determined manually as a piror [19, 10, 18, 17] in coordination with penalty factor without a common principle. Besides, VMD procedure works in complex field and to our knowledge there is no rigorous analysis on the convergence[5]. In order to alleviate the problem, the subsequent research either needs the knowledge to the range of the number of modes [11], or retrieves the IMFs recursively [12], or highly dependent on a predefined window-size to retrieve the peak and valley of spectra [6]. To our best knowledge, although there are some research focusing on wide-band extension of VMD [22], yet even to narrow-band signals, there are almost no complete research that can simultaneously be rigorous in theoretical analysis, while obtaining estimation of the number of modes and center frequency without any prior knowledge to the signal, and achieving non-recursive iterations.
In this article, we propose a variation-optimization based, globally convergent method to automatically determine the IMFs numbers and the corresponding center frequencies for narrow-banded signals. In short, we model the modes retrieving problem as adaptively extracting the supporting baseline of the signal spectrum amplitude function. By subtracting the extracted baseline from the signal spectrum amplitude function, we retrieve the significant frequency band components of the signal and extract the various modes of the narrow-band signal. The extracted supporting baseline is acquired by adversarially maximizing the integral of the baseline while minimizing the integral of its curvature. The supporting baseline is computed by iteratively solving a variational problem that is degenerated into the solution of a normal ordinary differential equation, and under proper condition and weight parameters the method always converges to the optimal curve that can be used to determine the center frequencies and the number of IMFs after proper estimation of the gap residual at the bottom. Once the number of IMFs and initialized frequencies are determined by our method, these parameters can be used to start the following VMD procedure. Experiments show that our method can quickly get accurate estimation about the number of IMFs and the corresponding initial center frequencies as prior to the following VMD procedure.
Our contribution can be summarized as 3 folds. First, as far as we know, in determining the number and central frequencies of IMFs, we are the first one to not only give a feasible algorithm that can decompose a narrow-banded signal automatically but also prove the global convergence in a rigorous mathematical view. Second, our proof generalizes the classical dual ascent method into functionals, which expands the scope of the application for the classical algorithm. Third, experiments show that our algorithm can filter out all the potential modes without any prior setting or assumptions to the number of modes, and also show good performance in both precision and speed even when the center frequencies of different modes are very close.
@@
2 Methodology
In this section we introduce our method for estimating the number of IMF modes as well as center frequencies of each band. We first introduce the background of the problem and give a brief introduction about Variational Mode Decomposition, and then we propose the concept of Supporting Baseline of a Function, finally we describe our method for finding the baseline and the whole workflow.
@@ 2.1 Supporting Baseline of a Function
Without loss of generality we first introduce the Supporting Baseline of a Function. For any function $s(x)>0$, the supporting baseline is a curve that tightly fits the lower bound of the function and reflects the overall trend of the lower bound. Fig. 1 gives an intuitive illustration about the conception of supporting baseline. As illustrated in Fig. 1, intuitively the supporting baseline should be tightly close to the lower bound of the function, approaching the infimum of functions as close as possible as well as maintain its own smoothness in order to capture the global trend of the lower bound rather than too much details, as capturing too much details will in turn reduce the robustness for separating the modes since there might be some noise added to the signal in either time or frequency domain.
@@ 2.2 Variational Methods for Finding the Supporting Baseline
To solve the problem for finding the supporting baseline, we use optimization method mainly based on variation and convex optimization [8, 4]. Let $f(x)\geq 0$ be a known function, the problem of finding the supporting baseline curve of $f(x)$ is equivalent to finding a function $g(x)$ that tightly approaching the infimum of $f(x)$ while maintaining the global trend of its lower bound. This can be viewed as the following optimization problem:
@@
then the above Lagrange Slack Function is like the form $J[g(x)]=\int_{\Omega}F(x,g,g^{\prime\prime})\text{d}x$ so that if we assume the variation is zero at both ends of the curve, one can easily know the optimal solution of the above function happen to satisfy the simplified form of Euler-Poisson equation [8] if we let the variation vanish, so it can be written as
@@
where 0 stands for the left end and $n$ the right end of the curve. To solve the boundary value problem of ODE, we use Finite Difference Method [20]. Note that for a differential function $g(x)$, it’s derivative of first to fourth order can be expressed approximately as the linear combination of $g(x)$ itself and we can take similar operation to the function $\alpha(x)$ (See Supplementary Materials 2). Since now different order of derivative with respect to $\alpha(x)$ and $g(x)$ can be expressed as the linear combination of themselves, one can easily transform (11) into the matrix-form as follows:
@@
where $\boldsymbol{A}^{(n)}$ and $\boldsymbol{G}^{(n)}$ is the conversion matrices (See Supplementary Materials 2) that transform the nth derivatives of $\alpha(x)$ and $g(x)$ into linear expression of the corresponding function, together with the boundary condition to form full-rank matrices, and $\boldsymbol{\alpha}$, $\boldsymbol{g}$, $\boldsymbol{\beta}$, $\boldsymbol{\lambda}$, $\boldsymbol{\mu}$ is obtained by respectively discretizing the corresponding functions, $\odot$ stands for component-wise multiplication between matrices(See Supplementary Materials 3 for details). The conversion matrices $\boldsymbol{A}^{(n)}$ and $\boldsymbol{G}^{(n)}$ share the same structure, in which the last four lines depicts the boundary condition and the other the ODE equation, and $\boldsymbol{g}$, $\boldsymbol{\alpha}$, $\boldsymbol{\beta}$, $\boldsymbol{\lambda}$, $\boldsymbol{\mu}$ are column vectors in which each element is a uniformly discrete sampled value on the function $g(x)$, $\alpha(x)$, $\beta(x)$, $\lambda(x)$, $\mu(x)$, respectively.
2.3 Proof of the Convergence
Algorithm 1 The Projection Gradient Descent Method for Evaluating Support Baseline
@@ 2.3 Proof of the Convergence
In this section we show the global convergence of the Algorithm 1. In our algorithm we have in fact solved a functional Extreme Value Problem with respect to the spectrum to the original signal. Despite the fact that the Fourier transform spectrum of a general signal may not necessarily be smooth, yet on one side, as in practical digital signal processing, the signal are discretized and can be smoothly interpolated, indicating the function processed by our algorithm are practically smooth which enable the reasonable use of commonly used assumptions that are advantageous for proving algorithm convergence, and on the other hand, subsequent numerical experiments have also shown that such assumptions not only do not affect the range of signals that can be processed, but are also very helpful for proving convergence. Therefore, in our proof, we assume that the spectrum is smooth so that we can extend the Lagrangian dual gradient ascent to functionals. Our proof combines functional, variational optimization and ODE theories in the framework of gradient projection, and can be divided into 3 steps. First we prove that we are dealing with a convex functional optimization problem, where the feasible domain is also a convex set. Next we show that under the condition of our algorithm, the Lagrangian dual gap is zero [4, 2] so that it is possible to find the optimum by dual ascent method with some modification from vectors to functional. After that, the convergence of iteration in dual space is shown and it is also shown that the optimal function can also be approached during the iterations by utilizing the dependence of differential equation solutions on coefficients. For the sake of limited spaces, we move the proof of convexity and duality into Supplementary Materials 4 and only give the proof to the convergence of iteration in Algorithm 1 in this part.
@@
We first restate the iteration in Algorithm 1. Notice that during the iteration, we have
@@
By using Lemma 15, we can prove that in iteration (14) the multiplier $\lambda_{k}(x)$, $\mu_{k}(x)$ always terminate at one of the optimum. To say in detail, we should have the following theorem:
@@
$(\lambda_{k+1}(x),\mu_{k+1}(x))$. If $\mathcal{G}$ is second-order differentiable and satisfies (17) so that $D^{2}\mathcal{G}$ is bounded with [L,L][-L,L]. Due to the concavity of $\mathcal{G}$ and the left side of the bound, and combine (30) we get
@@
if $\eta=\max\{\theta,\gamma\}=\cfrac{1}{L}<\cfrac{2}{L}$, then $\mathcal{G}$ is a monotonously increasing functional series with upper bound due to its concavity, so that $\mathcal{G}$ converges and the difference between adjacent term of both $\mathcal{G}$ and $(\lambda_{k}(x),\mu_{k}(x))$ is approaching zero.
Now we prove $\{g_{k}(x)\}$ is guaranteed to approach one of the optimum when the algorithm terminates even if $\mathcal{G}$ is not strongly concave. Due to the concavity the iteration on $\mathcal{G}$ should satisfy [2]
@@
(35) $+\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}$
since $\mathcal{G}$ is convex. Note that the spectrum of $D^{2}\mathcal{G}$ is larger than L-L and $\eta=\cfrac{1}{L}$ to replace the left side of (35), and use the concavity of $\mathcal{G}$, we have
$\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}\leq 2\eta(\mathcal{G}(\lambda_{k+1}(x),\mu_{k+1}(x))$
(36) $-\mathcal{G}(\lambda^{*}(x),\mu^{*}(x)))+\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k}(x)\\ \mu^{*}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
@@
$\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}\leq 2\eta\sum\limits_{i=0}^{k}(\mathcal{G}(\lambda_{i+1}(x),\mu_{i+1}(x))$
(37) $-\mathcal{G}(\lambda^{*}(x),\mu^{*}(x)))+\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{0}(x)\\ \mu^{*}(x)-\mu_{0}(x)\end{bmatrix}\right\|_{2}^{2}$

### v2

@@ Abstract
Automatically determining the number of intrinsic mode functions (IMFs) and their center frequencies in Variational Mode Decomposition (VMD) still remains an open mathematical challenge, which often relies on heuristic prior settings, trial-and-error strategies or complex-field based methods that lack of theoretical guarantees on the convergence. In this article, we proposed a novel variational method based on real field, focusing on evaluating the supporting baseline of a signal’s spectrum, to further retrieve the intrinsic mode functions. Our method treats automatic extraction of modes as a constrained optimization problem that adversarially maximizes the baseline integral while penalizing its curvature, and transform the problem into iteratively solving a fourth-order boundary value problem via Lagrangian duality. Furthermore, we establish the rigorous mathematical proof on the global convergence to our algorithm based on dual ascent in function space. Comprehensive numerical experiments on artificial and real-world signals including electrocardiogram (ECG) data show that our method can provide accurate estimates of IMFs and center frequencies, and comparison with methods like Successive VMD also shows our advantages in avoiding redundant modes while ensuring the recovery of necessary components, indicating that we have provided a robust, theoretically grounded initialization routine for VMD.
@@
narrow-banded signals;variational optimization;Intrinsic Mode Function
@@ 1 Introduction
Variational Mode Decomposition (VMD) decomposes original signals into multiple Intrinsic Mode Functions (IMF) with limited bandwidth with the help of variational optimization.[5] The core of VMD is to minimize the sum of the bandwidths with respect to the estimated IMFs as well as to ensure the sum of all the IMFs is equal to the original signal. In VMD, it is assumed the frequency of each mode is almost compact around a center pulsation and the associated analytic signal is computed by means of Hilbert Transform to get the unilateral frequency spectrum. Then each mode’s frequency spectrum was shifted into baseband by multiplying a frequency modulation factor with respect to the estimated center frequency, transforming each band-pass mode into low pass, and the bandwidth can be estimated by $H^{1}$ Gaussian smoothness. Afterwards, they decompose the original signal into each modes by minimizing the sum of the bandwidth of each mode with the constraints that the summation of all modes is the original signal. In short, their key variation optimization can be listed as following:
(1) $\min\limits_{u_{k},\omega_{k}}\left\{\sum\limits_{k}\left\|\partial_{t}\left[\left(\delta(t)+\frac{\text{j}}{\pi t}\right)*u_{k}(t)\right]e^{-\text{j}\omega_{k}t}\right\|_{2}^{2}\right\}$
(2) $\text{s.t.}\quad\sum\limits_{k}u_{k}=f$
so that each mode can be altered into an associated analytical signal by using Hilbert Transform and shifted into base-band for the following estimation of bandwidth using $H^{1}$ Gaussian smoothness with the shifted demodulated components. To solve the problem, one can construct an augmented Lagrangian Function by introducing both quadratic penalty and Lagrangian Multipliers as follows:
(3) $\mathcal{L}(u_{k},\omega_{k},\lambda)$ $=\alpha\sum_{k}\left\|\partial_{t}\left[\left(\delta(t)+\frac{j}{\pi t}\right)*u_{k}(t)\right]e^{-j\omega_{k}t}\right\|_{2}^{2}$
(4) $u_{k}^{n+1}=\underset{u_{k}\in\mathbb{R}}{\arg\min}\left\{\alpha\left\|\partial_{t}\left[\left(\delta(t)+\frac{j}{\pi t}\right)*u_{k}(t)\right]e^{-j\omega_{k}t}\right\|_{2}^{2}+\left\|f-\sum u_{i}+\frac{\lambda}{2}\right\|_{2}^{2}\right\}.$
(5) $\hat{u}_{k}^{n+1}=\underset{\hat{u}_{k},\,\hat{u}_{k}=\hat{u}_{k}^{T}}{\arg\min}\left\{\alpha\left\|j\omega\left[1+\text{sgn}(\omega+\omega_{k})\right]\hat{u}_{k}(\omega+\omega_{k})\right\|_{2}^{2}+\left\|\hat{f}-\sum\hat{u}_{i}+\frac{\lambda}{2}\right\|_{2}^{2}\right\}.$
(6) $\hat{u}_{k}^{n+1}=\underset{u_{k},\,\hat{u}_{k}=\hat{u}_{k}^{T}}{\arg\min}\left\{\int_{0}^{\infty}4\alpha(\omega-\omega_{k})^{2}|\hat{u}_{k}(\omega)|^{2}d\omega+2\int\left(\hat{f}-\sum\hat{u}_{i}+\frac{\hat{\lambda}}{2}\right)^{2}d\omega\right\}$
which can be evaluated by vanishing the first ordered variation of positive frequencies for both signal components and center frequencies
(7) $\hat{u}_{k}^{n+1}=\left(\hat{f}-\sum_{i\neq k}\hat{u}_{i}+\frac{\hat{\lambda}}{2}\right)\frac{1}{1+2\alpha(\omega-\omega_{k})^{2}}$
(8) $\omega_{k}^{n+1}=\frac{\displaystyle\int_{0}^{\infty}\omega\left|\hat{u}_{k}(\omega)\right|^{2}d\omega}{\displaystyle\int_{0}^{\infty}\left|\hat{u}_{k}(\omega)\right|^{2}d\omega}$
Since its establishment, VMD is always playing an important role in modern signal decomposition in many fields[14, 21, 9, 7, 16, 15, 1]. Although VMD has provided flexibility in decomposing signals, there are factors that hinder its further applications in that the number of intrinsic mode function must be determined manually as a prior in coordination with penalty factor without a common principle[19, 10, 18, 17]. In order to alleviate the problem, the subsequent research either needs the knowledge to the range of the number of modes ,[11] retrieves the IMFs recursively ,[12] or highly dependent on a predefined window-size to retrieve the peak and valley of spectra [6]. All these methods either require prior or rely on schemes devoid of convergence guarantees, making the results contains spurious or missed modes. The absence of a well-posed, convergent paradigm for automatically finding the number of IMFs is a crucial factor that hinders the development of VMD.
To close this gap, we introduce a fundamentally different approach for finding the IMFs. Instead of retrieving modes in an trial-and-error way or recursively extracting modes, we focus on the spectrum amplitude and reformulate the problem by introducing supporting baseline, a lower-enveloping curve of a signal’s spectrum which can support the spectrum as tightly as possible while keeping itself as smooth as possible, and transform the problem for finding IMFs into first evaluating the supporting baseline by adversarially maximizing its integral while penalizing its curvature, then retrieving the separated peaks of the original spectrum above the supporting baseline.
Our main contribution can be summarized as 4 folds. First, we give a new perspective for determining the number of IMFs by introducing supporting baseline of spectrum and solve the problem in real field. Second, we derive its optimality conditions, establishing a rigorous equivalence between this variational problem and a fourth-order boundary value problem, and establish rigorous mathematical proof for the global convergence to the course of dual ascent iteration in function space. Thirdly, we develop an efficient iteration scheme for the algorithm by converting the corresponding difference equation and boundary conditions into matrix format and expanding compatible rule to the component-wise multiplication between matrices and vectors for Hadamard product. Finally, we give experiments on both artificial and real signals, and comparison with SVMD to show the robustness of our algorithm and experiments show that our method can quickly get accurate estimation about the number of IMFs and the corresponding initial center frequencies as prior to the following VMD procedure.
@@
2 Methodology and Convergence Proof
In this section we introduce our method for estimating the number of IMF modes as well as center frequencies of each band. We first introduce the background of the problem, and then we propose the concept of Supporting Baseline of a Function, finally we describe our method for finding the baseline and give the convergence proof for our method.
@@ 2.1 Supporting Baseline of a Function
Without loss of generality we first introduce the Supporting Baseline of a function. For any function $f(x)>0$, the supporting baseline is a curve that tightly fits the lower bound of the function while reflects the overall trend of the lower bound. Fig. 1 gives an intuitive illustration about the conception of supporting baseline. As illustrated in Fig. 1, intuitively the supporting baseline should be tightly close to the lower bound of the function, approaching the infimum of functions as close as possible as well as maintain its own smoothness in order to capture the global trend of the lower bound rather than too much details, as capturing too much details will in turn reduce the robustness for separating the modes since there might be some noise added to the signal in either time or frequency domain.
@@ 2.2 Variational Methods for Finding the Supporting Baseline
To solve the problem for finding the supporting baseline, we use optimization method mainly based on variation and convex optimization[8, 4]. Let $f(x)\geq 0$ be a known function, the problem of finding the supporting baseline curve of $f(x)$ is equivalent to finding a function $g(x)$ that tightly approaching the point-wise infimum of $f(x)$ while maintaining the global trend of its lower bound. This can be viewed as the following optimization problem:
@@
then the above Lagrange Slack Function is like the form $J[g(x)]=\int_{\Omega}F(x,g,g^{\prime\prime})\text{d}x$ so that if we assume the variation is zero at both ends of the curve, one can easily know the optimal solution of the above function happen to satisfy the simplified form of Euler-Poisson equation if we let the variation vanish,[8] so it can be written as
@@
where 0 stands for the left end and $n$ the right end of the curve. To solve the boundary value problem of ODE, we use Finite Difference Method.[20] Note that for a differential function $g(x)$, it’s derivative of first to fourth order can be expressed as the linear combination of $g(x)$ itself in the corresponding differential format and we can take similar operation to the function $\alpha(x)$ (See Supplementary Materials A), so that different orders of derivative with respect to $\alpha(x)$ and $g(x)$ can be expressed as the linear combination of themselves, one can easily transform (19) into the matrix-form as follows:
@@
[…]
Algorithm 1 The Key Steps for Evaluating Support Baseline
@@
2.3 Proof of the Convergence
One of the core contribution of this work is to provide a deterministic, global convergence guarantees for the problem of automatic mode detection, which is typically unaddressed in the pioneer works. we prove that the algorithm generates sequences converging to an optimal solution of the original problem. The proof is built upon several folds:
Convexity and Duality: Proof of the problem’s convexity and strong duality.
Well-posedness of Iterations: Demonstration that each iterative step solves a well-posed subproblem with a unique solution.
Convergence in Dual Space: Impose functional functional to dual ascent method and analyse its convergence under both unique and non-unique optimal dual solutions.
Convergence in Primal Space: Establish the convergence of the primal sequence via the continuous dependence of solutions on parameters.
For the conciseness of the proof, we leave the detail of the the convexity and duality into Appendix and focus on the well-posedness of iterations and convergence in dual and primal spaces here.
2.3.1 Convexity and Duality
In this part we state the main results about the convexity and duality, and we put the details in Appendix B.1 and B.2.
Lemma 2.1.
The objection functional of (10) is convex, and the feasible set is also convex.
Lemma 2.2.
The dual of objection functional of (10) is concave, with the dual-gap zero.
2.3.2 Well-posedness of Iterations
Before analyzing the convergence of the iterative algorithm, we must ensure that the subproblem solved at each iteration is well-posed. Specifically, as in each iteration, our primal solution $g_{k+1}(x)$ is given by dual variables $(\lambda_{k}(x),\mu_{k}(x))$ at the k-th step and we transform the variational problem into a linear fourth-order boundary value problem (BVP), the well-posedness of this BVP encompassing the existence and uniqueness of the solution is crucial to guarantee that our iterative scheme is not only well-defined but also stable against small numerical perturbations. Since $\alpha(x)$ and $\beta(x)$ in our algorithm can be considered as weights function that are always positive, we can establish the existence and uniqueness of our solution to (19) below.
Theorem 2.3.
Assume $\alpha(x)>0$, then the solution of (19) with $g(0)=0,\\ g(n)=0,g^{\prime}(0)=0,g^{\prime}(n)=0$ is $g(x)\equiv 0$.
Proof 2.4.
Recall that when dealing with the homogeneous condition, we exactly have
where
(23) $g(a)=g(b)=g^{\prime}(a)=g^{\prime}(b)=0$
It is obvious that $C_{3}$ and $C_{4}$ are zero by $g(a)$ and $g^{\prime}(a)$, and $g^{\prime}(b)=0$ is equivalent to
(24) $\int_{a}^{b}\cfrac{C_{1}t+C_{2}}{\alpha(t)}\text{d}t=0$
Since $\alpha(t)>0$ and $\alpha(t)\in C[a,b]$, since $C_{1}t+C_{2}$ is a linear function so that if the zero point is not in $[a,b]$ then it is not possible to make (24) hold. If there exists $t\in[a,b]$ that makes $C_{1}t+C_{2}$ zero, then the upper limit integral function $\Phi(x)=\displaystyle\int_{a}^{x}\cfrac{C_{1}t+C_{2}}{\alpha(t)}\text{d}t$ has no zero points in $(a,b)$, thus makes $g(b)=\displaystyle\int_{a}^{b}\Phi(t)\text{d}t$ non-zero, which is paradoxical to our assumption if not both $C_{1}$ and $C_{2}$ are zero. Then according to Fredholm Alternative Theorem, we can assert that during each iteration, there always exist a unique solution to g(x)g(x)[26, 27, 28].
Corollary 2.5.
For each iteration of (21), there always exists a unique solution.
2.3.3 Convergence in Dual Space
To facilitate the convergence analysis in the dual space, we introduce the following regularity assumptions on the primal problem:
The target spectrum function $f(x)$ is Lipschitz continuous on $\mathcal{D}$.
The solution $g(x)$ to the subproblem and the iterates $g_{k}(x)$ reside in $C^{2}(\mathcal{D})$ with uniformly bounded second derivatives.
These assumptions are standard for analyzing variational problems involving curvature penalties. In practice, for discrete and potentially noisy spectral data, $f(x)$ can be considered as a smooth interpolation of the observed points. Crucially, the core object of interest—the supporting baseline—is intrinsically a smooth function due to the curvature penalty term in the objective. Therefore, the analysis under these smoothness assumptions captures the essential behavior of the algorithm.
@@ 2.3.3 Convergence in Dual Space
We begin the proof by first restating the iteration in Algorithm 1. Notice that during the iteration, we have
@@
Now we prove that in iteration (25) the multiplier $\lambda_{k}(x)$, $\mu_{k}(x)$ always terminate at one of the optimum. To say in detail, we should have the following theorem:
@@
$(\lambda_{k+1}(x),\mu_{k+1}(x))$. If $\mathcal{G}$ is second-order differentiable and satisfies (2.8) so that $D^{2}\mathcal{G}||$ is bounded with [L,L][-L,L]. Due to the concavity of $\mathcal{G}$ and the left side of the bound, and combine (41) we get
@@
if $\eta=\max\{\theta,\gamma\}=\frac{1}{L}<\frac{2}{L}$, then $\mathcal{G}$ is a monotonously increasing functional series with upper bound due to its concavity, so that $\mathcal{G}$ converges and the difference between adjacent term of both $\mathcal{G}$ and $(\lambda_{k}(x),\mu_{k}(x))$ is approaching zero.
Now we prove $\{g_{k}(x)\}$ is guaranteed to approach one of the optimum when the algorithm terminates even if $\mathcal{G}$ is not strongly concave. By introducing $l(\boldsymbol{y};\boldsymbol{x})=\mathcal{G}(\boldsymbol{x})+D\mathcal{G}(\boldsymbol{x})(\boldsymbol{y}-\boldsymbol{x})$ and due to the concavity the iteration on $\mathcal{G}$ should satisfy [2]
@@
(46) $+\cfrac{1}{2\eta}\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}$
since $\mathcal{G}$ is convex. Due to the concavity of $\mathcal{G}$, we have
(47) $\mathcal{G}(\lambda_{k+1}(x),\mu_{k+1}(x))\geq l\left(\begin{bmatrix}\lambda_{k+1}(x)\\ \mu_{k+1}(x)\end{bmatrix};\begin{bmatrix}\lambda_{k}(x)\\ \mu_{k}(x)\end{bmatrix}\right)-\cfrac{1}{2\eta}\left\|\begin{bmatrix}\lambda_{k+1}(x)-\lambda_{k}(x)\\ \mu_{k+1}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
and
(48) $l\left(\begin{bmatrix}\lambda^{*}(x)\\ \mu^{*}(x)\end{bmatrix};\begin{bmatrix}\lambda_{k}(x)\\ \mu_{k}(x)\end{bmatrix}\right)\geq\mathcal{G}(\lambda^{*}(x),\mu^{*}(x))$
Substituting (47) and (48) into (46), then we have
$\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}$ $\leq 2\eta[\mathcal{G}(\lambda_{k+1}(x),\mu_{k+1}(x))$
(49) $-\mathcal{G}(\lambda^{*}(x),\mu^{*}(x))]+\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k}(x)\\ \mu^{*}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
@@
$\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}$ $\leq 2\eta\sum\limits_{i=0}^{k}[\mathcal{G}(\lambda_{i+1}(x),\mu_{i+1}(x))$
(50) $-\mathcal{G}(\lambda^{*}(x),\mu^{*}(x))]+\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{0}(x)\\ \mu^{*}(x)-\mu_{0}(x)\end{bmatrix}\right\|_{2}^{2}$